I need to see how to work this problem:
On Friday Clothiers sold some jeans at $25 and shirts at $18 Receipts total $441
Sat. both items priced at $20 each sold exactly the same number of each item and had receipts of $420, how many pairs of jeans and how many shirts were sold each day?

Answer by lwsshak3(11628)

On Friday Clothiers sold some jeans at $25 and shirts at $18 Receipts total $441
Sat. both items priced at $20 each sold exactly the same number of each item and had receipts of $420, how many pairs of jeans and how many shirts were sold each day?
**
let x=no. of jeans sold each day
let y=no. of shirts sold each day
..
25x+18y=441 (Friday)
20x+20y=420 (Saturday)
..
multiply 1st equation by 4:
100x+72y=1764
multiply 2nd equation by 5:
100x+100y=2100
subtract 2nd equation from 1st equation
-28y=-336
y=12
20x=420-20y=420-240=180
x=180/20=9
no. of jeans sold each day=9
no. of shirts sold each day=12